# Valentina Bais and Rafael Torres Scuola Internazionale Superiori di Studi Avanzati (SISSA) Via Bonomea 265 34136

### v2

@@
Abstract: We introduce a simple cut-and-paste mechanism to construct both orientable and nonorientable four-manifolds from a given initial one. This mechanism alters the fundamental group while preserving other essential topological invariants. It avoids codimension two cut-and-paste fundamental group computations and fast tracks the search for fixed-point free involutions. The mechanism proves useful to unveil novel exotic irreducible smooth structures on closed four-manifolds with finite cyclic fundamental group, which include $\mathbb{Q}$-homology real projective four-spaces.
@@ 3. III
The crux to the efficiency of the procedure is its fourth step. In principle, one would hope that the diffeomorphism invariant that discerns $M_{1}$ from $M_{2}$ also tells apart the new four-manifolds $M_{1}(p)$ and $M_{2}(p)$. This is indeed the case in all the examples under consideration in this paper, where we reckoned the Heegaard Floer homology based invariants used by Levine-Lidman-Piccirillo in [29], the Ozsváth-Szabó invariants [31], the Seiberg-Witten invariants [20], [34, Section 2.4] (as employed by Stipsicz-Szabó [33, 34] and Baykur-Stipsicz-Szabó [8]), and the $\eta$-invariant as studied by Stolz [35].
As a warm up result, we paint a picture of the fourth step and the utility of the mechanism through an extension of a recent result of Levine-Lidman-Piccirillo. In [29], the authors constructed the first example of an inequivalent smooth structure on a closed smooth oriented four-manifold with definite intersection form: they built a four-manifold $\mathcal{R}$ that is homeomorphic but not diffeomorphic to the connected sum $\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$, where $\Sigma_{2}$ is a $\mathbb{Q}$-homology four-sphere with $\pi_{1}(\Sigma_{2})=\mathbb{Z}/2$. Using the pair $(\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}},\mathcal{R})$ in the first step of our mechanism, we obtain the following generalization of their result.
@@
For every odd $p\in\mathbb{Z}_{>0}$, there is a closed smooth oriented four-manifold $\mathcal{R}_{2p}$ that is homeomorphic but not diffeomorphic to the connected sum $\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$, where $\Sigma_{2p}$ is a $\mathbb{Q}$-homology four-sphere with cyclic fundamental group of order $2p$. The diffeomorphism types are distinsguished by the Ozsváth-Szabó invariant.
To distinguish the smooth structures in [29, Theorem 1.2], Levine-Lidman-Piccirillo show that the universal covers ${\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{E}$ of the four-manifolds $\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{R}$ are non-diffeomorphic by means of their Ozsváth-Szabó invariant [31]. For every odd $p\in\mathbb{Z}_{>0}$, we show that the degree-two coverings of our four-manifolds are obtained from ${\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{E}$, respectively, by taking a connected sum with a $\mathbb{Q}$-homology four-sphere. This observation, together with the properties of the Heegaard Floer smooth four-dimensional invariant, allows us to canonically adapt the proof of [29, Theorem 1.2] to distinguish such coverings and, hence, also the four-manifolds $\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{R}_{2p}$.
Our next theorem was inspired by recent work of Stipsicz-Szabó [33, 34] and Baykur-Stipsicz-Szabó [8], where they built infinitely many nondiffeomorphic irreducible smooth structures on $\Sigma_{4m+2}\mathbin{\#}\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ for $\Sigma_{4m+2}$ a $\mathbb{Q}$-homology four-sphere with fundamental group $\mathbb{Z}/(4m+2)$. Recall that a smooth four-manifold $M$ is said to be irreducible if for every smooth connected sum decomposition $M=M_{1}\mathbin{\#}M_{2}$ either $M_{1}$ or $M_{2}$ is a homotopy four-sphere [20, Definition 10.1.17].
@@
  • The four-manifold $\widetilde{X}_{n}$ has exactly two Seiberg-Witten basic classes ±$\pm K_{n}\in\mathcal{B}_{\widetilde{X}_{n}}$ and $SW_{\widetilde{X}_{n}}(\pm K_{n})=\pm n^{2}$.
@@
If the four-manifold $\widetilde{X}_{n}$ does admit an involution (1.1), but has more than two Seiberg-Witten basic classes, assume it is minimal. Then, there exists a closed smooth oriented irreducible four-manifold $X_{n}(2p)$ with finite fundamental group of order $2p$ for which there is a homeomorphism (1.2) whose double cover is (1.3).
In [33, 34, 8], the Seiberg-Witten invariants are used to distinguish the diffeomorphism types of their examples in the following way. Notice that a diffeomorphism $h:M_{1}\rightarrow M_{2}$ between two closed oriented four-manifolds lifts to a diffeomorphism of the degree-two coverings $\hat{h}:\widehat{M}_{1}\rightarrow\widehat{M}_{2}$, whenever the index-two subgroups of $\pi_{1}(M_{i})$ identifying $\widehat{M}_{i}\rightarrow M_{i}$ correspond under the map induced by $h$; see Lemma 9. Stipsicz-Szabó obstruct the existence of the diffeomorphism $h$ through the diffeomorphism invariance of the Seiberg-Witten function $\operatorname{SW}_{\widehat{M}_{i}}:\operatorname{Spin}^{\mathbb{C}}(\widehat{M}_{i})\rightarrow\mathbb{Z}$ for $i=1,2$. Moreover, they conclude on the irreducibility of $\widehat{M}_{i}$ and $M_{i}$ through their Seiberg-Witten basic classes. We follow the same approach in the proof of Theorem D. In our case, however, the degree-two coverings of the four-manifolds of Theorem D do decompose smoothly as a connected sum $M\mathbin{\#}\Sigma_{p}$. In the proof of Theorem D, we couple the computations of the Seiberg-Witten invariants of Stipsicz-Szabó with a result of Kotschick-Morgan-Taubes [27, Proposition 2] that states that there are $\text{Spin}^{\mathbb{C}}$-structures on $M$ and $M\mathbin{\#}\Sigma_{p}$ for which the values of their Seiberg-Witten invariants coincide; see Section 4.1.
@@
$m{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\#n\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}\#\Sigma_{2p}.$
@@
Concerning the nonorientable realm, we construct new smooth structures on $\mathbb{Q}$-homology real projective four-spaces and non-smoothable copies of these four-manifolds through our mechanism, and extend results of Cappell-Shaneson [10], Fintushel-Stern [12] and Ruberman [32]. In order to build the novel four-manifolds, we start with Cappell-Shaneson’s exotic $\mathbb{R}P^{4}$ or the non-smoothable four-manifold $\ast\mathbb{R}P^{4}$ that is homotopy equivalent to the real projective four-space. We address the fourth step of our construction mechanism by computing the $\eta$-invariant [35] of the smooth four-manifolds involved in Section 6.4. Our main result in this case is the following theorem.
@@
    The connected sums $A(2p)\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}$ and $B(2p)\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}$ are diffeomorphic.
Moreover, there is a non-smoothable closed topological four-manifold $\ast A(2p)$ that is homotopy equivalent to $A(2p)$.
We have collected quintessential properties of the known inequivalent smooth structures on nonorientable four-manifolds in the statement of Theorem F for the convenience of the reader.
@@ 1.1. Acknowledgements
V. B. is a member of GNSAGA – Istituto Nazionale di Alta Matematica ‘Francesco Severi’, Italy. R. T. thanks Stefan Bauer, Bob Gompf, Paul Kirk, Tye Lidman, Porter Morgan, Tom Mrowka, Mark Powell, Cliff Taubes, Mike Usher and Zoltan Szabó for useful e-mail exchanges.
@@ 2.1. Cut-and-paste procedure: oriented setting
We construct the compact oriented four-manifold $S_{p}$ of the second step of our procedure as follows. Let $\alpha\subset S^{1}\times S^{3}$ be a simple loop whose homotopy class generates the infinite cyclic group $\pi_{1}(S^{1}\times S^{3})=\mathbb{Z}$. Denote by $\alpha_{p}\subset S^{1}\times S^{3}$ the simple loop that represents the homology class $[\alpha]^{p}\in H_{1}(S^{1}\times S^{3};\mathbb{Z})=\mathbb{Z}$. The tubular neighborhood of this codimension three submanifold is $\nu(\alpha_{p})=S^{1}\times D^{3}$. The main building block for our cut-and-paste construction in the orientable setting is the compact four-manifold
@@
The closed smooth oriented four-manifold
@@
A description of the cyclic coverings of the building block $S_{p}$ is given in our next lemma, which is of great utility in the sequel.
@@
$\bullet$ For every $p\in\mathbb{Z}_{>0}$, there is a double covering
@@
where $\{B_{1},\dots,B_{p}\}\subset S^{3}$ are disjointly embedded smooth $3$-balls. The $\mathbb{Z}$-action on $\widetilde{S}_{p}$ is generated by the map
$\phi:\widetilde{S}_{p}\rightarrow\widetilde{S}_{p}$
$(t,x)\mapsto(t+1,f_{p}(x))$
where $f_{p}$ is the restriction of a smooth automorphism of $S^{3}$ which cyclically permutes the $3$-balls $B_{1},\dots,B_{p}$. The degree-two covering of $S_{p}$ associated to the subgroup $2\mathbb{Z}\subset\mathbb{Z}=\pi_{1}(S_{p})$ is given by the quotient of $\widetilde{S}_{p}$ by the $\mathbb{Z}$-action generated by the map $\phi^{2}$. The conclusion follows from the fact that the square of a cyclic permutation of odd length is still cyclic, while the square of a cyclic permutation of even length factors as a product of two cycles of half the original length.
@@
In the following, it will be useful to consider the genus one Heegaard splitting of the 3-sphere
@@
We can assume that the map $f_{p}$ in the proof of Lemma 2 is a rotation of $S^{3}$ of angle $2\pi/p$, which restricts to a rotation along the $D^{2}$-factor on the solid torus on the left side of (2.3) and to a rotation along the $S^{1}$-factor on the solid torus on the right side. Moreover, we can choose the three-balls $B_{1},\dots,B_{p}$ to be symmetrically embedded in the solid torus on the right side of (2.3), where they are permuted by the action of $f_{p}$.
@@
[Proof]A result of Gluck [17] states that the diffeotopy group of $S^{1}\times S^{2}$ is isomorphic to $\mathbb{Z}/2$ and it is generated by the self-diffeomorphism
@@
given by
$(\theta,x)\mapsto(\theta,r_{\theta}(x)),$
where $r_{\theta}$ is the rotation of $S^{2}$ of angle $\theta$ around a fixed axis [17], [20, p. 156]. Such map is known as a Gluck twist. In order to prove Theorem 4, it is enough to show that the diffeomorphism $\varphi$ admits an extension $\Phi_{p}:S_{p}\rightarrow S_{p}$ for any odd $p\in\mathbb{Z}_{>0}$. We explicitly build the diffeomorphism $\Phi_{p}$ by using the identification
$S_{p}\cong(\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i}))/\sim_{\phi}$
coming from Lemma 2 as follows. Without loss of generality, we work in the setting of Remark 3. In particular, the deck transformation $\phi$ and the collection of three-balls $B_{1},\dots,B_{p}\subset D^{2}\times S^{1}$ are defined as in the proof of Lemma 2.
@@
$F:\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})\rightarrow S^{3}\setminus\sqcup_{i=1}^{p}B_{i}$
@@
where $r_{t}$ denotes the rotation of $S^{3}$ of an angle $2\pi t$ that restricts to a rotation along the $D^{2}$-factor on the solid torus $D^{2}\times S^{1}$ of the Heegaard decomposition of $S^{3}$ in Remark 3. Such solid torus contains the three-balls $B_{1},\dots,B_{p}$. Notice that for every $t\in[0,1]$, the map $F(t,\cdot)$ is an $f_{p}$-equivariant self-diffeomorphism of $S^{3}\setminus\sqcup_{i=1}^{p}B_{i}$, where $f_{p}$ can be chosen to act on the genus one Heegaard splitting of $S^{3}$ described in Remark 3. We can now define the automorphism
$\Phi:\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})\rightarrow\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})$
$(t,x)\mapsto(t,F(t,x)).$
It is immediate to check that $\Phi$ is $\phi$-equivariant and, thus, it induces a self-diffeomorphism $\Phi_{p}:S_{p}\rightarrow S_{p}$. In particular, the restriction of $\Phi_{p}$ to $\partial S_{p}=S^{1}\times S^{2}$ is a composition of $p$ Gluck twists, which is again a Gluck twist since $p$ is odd. This concludes the proof.
As mentioned in the first step of our introduction, remove the tubular neighborhood of a framed simple loop $\alpha_{M}\subset M$ from a closed oriented four-manifold $M$ to obtain a compact oriented four-manifold $M\setminus\nu(\alpha_{M})$ with boundary $\partial(M\setminus\nu(\alpha_{M}))=S^{1}\times S^{2}$. Cap off its boundary with the compact four-manifold defined in (2.1) to construct a closed oriented four-manifold
@@
The reader is directed to [16, Chapter 5] regarding existence and uniqueness results of the tubular neighborhood of a locally flat simple loop in a topological four-manifold.
@@
[Proof]We provide a proof of the third item. Let $\alpha_{M}\subset M$ be a simple loop as in (2.5) and denote by $\widetilde{\alpha}_{M}\subset\widetilde{M}$ its preimage in the universal cover $\widetilde{M}$ of $M$. The universal cover $\widetilde{M}(p)$ of $M(p)$ is constructed by gluing together copies of $\widetilde{M}\setminus\nu(\widetilde{\alpha}_{M})$ and $\widetilde{S}_{p}/\sim_{\phi^{k}}$ by matching each boundary component of $\widetilde{M}\setminus\nu(\widetilde{\alpha}_{M})$ with a boundary component of $\widetilde{S}_{p}/\sim_{\phi^{k}}$, where $\widetilde{S}_{p}$ is the universal covering of $S_{p}$ (see Lemma 2) and $k$ is the order of $[\alpha_{M}]\in\pi_{1}(M)$. In particular, if $k=\infty$ we set $\widetilde{S}_{p}/\sim_{\phi^{\infty}}=\widetilde{S}_{p}$. It follows that the equivariant intersection form $\widetilde{I}_{M(p)}$ of $M(p)$ is
@@
A proof of the other items of Lemma 6 is obtained from straight-forward homological computations, Novikov’s additivity and the Seifert-van Kampen theorem.
@@
Let $M$ be a closed oriented four-manifold and let $\alpha_{M}\subset M$ be a simple framed loop whose homotopy class is $[\alpha_{M}]=0\in\pi_{1}(M)$. There is a diffeomorphism

### v3

@@
Abstract: We introduce a simple cut-and-paste mechanism to construct both orientable and nonorientable four-manifolds from a given initial one. This mechanism alters the fundamental group while preserving other essential topological invariants. It avoids codimension two cut-and-paste fundamental group computations and fast tracks the search for fixed-point free involutions. The mechanism proves useful to unveil novel exotic irreducible smooth structures on closed four-manifolds with finite cyclic fundamental group, which include $\mathbb{Q}$-homology real projective four-spaces, as well as non-smoothable examples.
@@ 3. III
The crux to the efficiency of the procedure is its fourth step. In principle, one would hope that the diffeomorphism invariant that discerns $M_{1}$ from $M_{2}$ also tells apart the new four-manifolds $M_{1}(p)$ and $M_{2}(p)$. This is indeed the case in all the examples under consideration in this paper, where we reckoned the Heegaard Floer homology based invariants used by Levine-Lidman-Piccirillo in [32], the Ozsváth-Szabó invariants [35], the Seiberg-Witten invariants [21], [40, Section 2.4] (as employed by Stipsicz-Szabó [39, 40] and Baykur-Stipsicz-Szabó [8]), and the $\eta$-invariant as studied by Stolz [41].
As a warm up result, we paint a picture of the fourth step and the utility of the mechanism through an extension of a recent result of Levine-Lidman-Piccirillo. In [32], the authors constructed the first example of an inequivalent smooth structure on a closed smooth oriented four-manifold with definite intersection form: they built a four-manifold $\mathcal{R}$ that is homeomorphic but not diffeomorphic to the connected sum $\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$, where $\Sigma_{2}$ is a $\mathbb{Q}$-homology four-sphere with $\pi_{1}(\Sigma_{2})=\mathbb{Z}/2$. Using the pair $(\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}},\mathcal{R})$ in the first step of our mechanism, we obtain the following generalization of their result.
@@
For every odd $p\in\mathbb{Z}_{>0}$, there is a closed smooth oriented four-manifold $\mathcal{R}_{2p}$ that is homeomorphic but not diffeomorphic to the connected sum $\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$, where $\Sigma_{2p}$ is a $\mathbb{Q}$-homology four-sphere with cyclic fundamental group of order $2p$. The diffeomorphism types are distinsguished by the Ozsváth-Szabó invariant.
To distinguish the smooth structures in [32, Theorem 1.2], Levine-Lidman-Piccirillo show that the universal covers ${\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$ and $\mathcal{E}$ of the four-manifolds $\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$ and $\mathcal{R}$ are non-diffeomorphic by means of their Ozsváth-Szabó invariant [35]. For every odd $p\in\mathbb{Z}_{>0}$, we show that the degree-two covers of our four-manifolds are obtained from ${\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$ and $\mathcal{E}$, respectively, by taking a connected sum with a $\mathbb{Q}$-homology four-sphere. This observation, together with the properties of the Heegaard Floer smooth four-dimensional invariant, allows us to canonically adapt the proof of [32, Theorem 1.2] to distinguish such covers and, hence, also the four-manifolds $\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$ and $\mathcal{R}_{2p}$.
Our next theorem was inspired by recent work of Stipsicz-Szabó [39, 40] and Baykur-Stipsicz-Szabó [8], where they built infinitely many nondiffeomorphic irreducible smooth structures on $\Sigma_{4m+2}\mathbin{\#}\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$ for $\Sigma_{4m+2}$ a $\mathbb{Q}$-homology four-sphere with fundamental group $\mathbb{Z}/(4m+2)$. Recall that a smooth four-manifold $M$ is said to be irreducible if for every smooth connected sum decomposition $M=M_{1}\mathbin{\#}M_{2}$ either $M_{1}$ or $M_{2}$ is a homotopy four-sphere [21, Definition 10.1.17]. A smooth four-manifold is minimal if it contains no smoothly embedded two-sphere of self-intersection minus one. The presence of such two-sphere in a smooth four-manifold $M$ guarantees a smooth connected sum decomposition $M=M_{0}\mathbin{\#}\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$ for some four-manifold $M_{0}$ [21, Proposition 2.2.11].
@@
  • The four-manifold $\widetilde{X}_{n}$ has exactly two Seiberg-Witten basic classes $\pm K_{n}\in\mathcal{B}_{\widetilde{X}_{n}}$ and $SW_{\widetilde{X}_{n}}(\pm\mathfrak{s}_{n})=\pm n^{2}$ for $c_{1}(\pm\mathfrak{s}_{n})=\pm K$.
@@
If the four-manifold $\widetilde{X}_{n}$ does admit an involution (1.1), but has more than two Seiberg-Witten basic classes, assume it is minimal and symplectic with $b_{2}^{+}(\widetilde{X})>1$. Then, there exists a closed smooth oriented irreducible four-manifold $X_{n}(2p)$ with finite fundamental group of order $2p$ for which there is a homeomorphism (1.2) whose double cover is (1.3).
In [39, 40, 8], the Seiberg-Witten invariants are used to distinguish the diffeomorphism types of their examples in the following way. A diffeomorphism $h:M_{1}\rightarrow M_{2}$ between two closed oriented four-manifolds lifts to a diffeomorphism of the degree-two covers $\hat{h}:\widehat{M}_{1}\rightarrow\widehat{M}_{2}$, whenever the index-two subgroups of $\pi_{1}(M_{i})$ identifying $\widehat{M}_{i}\rightarrow M_{i}$ correspond under the map induced by $h$; see Lemma 9. Stipsicz-Szabó obstruct the existence of $h$ through the diffeomorphism invariance of the Seiberg-Witten function $\operatorname{SW}_{\widehat{M}_{i}}:\operatorname{Spin}^{\mathbb{C}}(\widehat{M}_{i})\rightarrow\mathbb{Z}$ for $i=1,2$. Moreover, they conclude on the irreducibility of $\widehat{M}_{i}$ and $M_{i}$ through their Seiberg-Witten basic classes. We follow the same approach to prove Theorem D. In our case, however, the degree-two covers do decompose smoothly as a connected sum $M\mathbin{\#}\Sigma_{p}$, and we build on arguments of Baykur-Stipsicz-Szabó [8] and Park [37]; see Section 4.1.
@@
$m{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}\#n\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}\#\Sigma_{2p}.$
@@
Concerning the nonorientable realm, we construct new smooth structures on $\mathbb{Q}$-homology real projective four-spaces and non-smoothable copies of these four-manifolds through our mechanism. This extends results of Cappell-Shaneson [10], Fintushel-Stern [13] and Ruberman [38]. In order to build the novel four-manifolds, we start with either Cappell-Shaneson’s exotic ${\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}$ or the non-smoothable four-manifold $\ast{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}$ that is homotopy equivalent to the real projective four-space; see [47, Chapter 14], [38, Section 2]. We address the fourth step of our construction mechanism by computing the $\eta$-invariant [41] of the smooth four-manifolds involved in Section 6.4. Our main result in this case is the following theorem.
@@
    The connected sums $A(2p)\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}$ and $B(2p)\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}$ are diffeomorphic.
Moreover, there is a non-smoothable closed topological four-manifold $\ast A(2p)$ that is homotopy equivalent to $A(2p)$ and whose universal cover is homeomorphic to $(p-1)(S^{2}\times S^{2})$.
Theorem F provides a pair of smooth orientation-reversing free $\mathbb{Z}/2p$-actions on a connected sum $(p-1)(S^{2}\times S^{2})$ that are equivariantly homeomorphic, but not equivariantly diffeomorphic. This extends work of Fintushel-Stern [13], and Cappell-Shaneson [10] and Gompf [20]. The last clause of Theorem F gives examples of topological orientation-reversing free $\mathbb{Z}/2p$-actions on $(p-1)(S^{2}\times S^{2})$ whose orbit space is non-smoothable; see Hambleton-Kreck-Teichner [24]. Moreover, we have collected quintessential properties of the known inequivalent smooth structures on nonorientable four-manifolds in the statement of Theorem F for the convenience of the reader.
@@ 1.1. Acknowledgements
We thank Stefan Bauer, Bob Gompf, Paul Kirk, Tye Lidman, Porter Morgan, Lisa Piccirillo and Zoltan Szabó for useful e-mail exchanges/conversations.
@@ 2.1. Cut-and-paste procedure: oriented setting
We construct the compact oriented four-manifold $S_{p}$ of the second step of our procedure as follows. Let $\alpha\subset S^{1}\times S^{3}$ be a simple loop whose homotopy class generates the infinite cyclic group $\pi_{1}(S^{1}\times S^{3})=\mathbb{Z}$. Denote by $\alpha_{p}\subset S^{1}\times S^{3}$ the simple loop that represents the homology class $\alpha]^{p}\in H_{1}(S^{1}\times S^{3};\mathbb{Z})=\mathbb{Z}$. The tubular neighborhood of this codimension three submanifold is $\nu(\alpha_{p})=S^{1}\times D^{3}$. The main building block for our cut-and-paste construction in the orientable setting is the compact four-manifold
@@
The closed oriented four-manifold
@@
A description of the cyclic covers of the building block $S_{p}$ is given in our next lemma, which is of great utility in the sequel.
@@
$\bullet$ For every $p\in\mathbb{Z}_{>0}$, there is a double cover
@@
where $\{B_{1},\dots,B_{p}\}\subset S^{3}$ are disjointly embedded smooth three-balls. The $\mathbb{Z}$-action on $\widetilde{S}_{p}$ is generated by the map
(2.3) $\phi:\widetilde{S}_{p}\rightarrow\widetilde{S}_{p}$
given by
(2.4) $(t,x)\mapsto(t+1,f_{p}(x)),$
where $f_{p}:(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})\rightarrow(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})$ is the restriction of a self-diffeomorphism of the three-sphere that cyclically permutes the three-balls $\{B_{1},\dots,B_{p}\}$. The degree-two cover of $S_{p}$ associated to the subgroup $2\mathbb{Z}\subset\mathbb{Z}=\pi_{1}(S_{p})$ is given by the quotient of $\widetilde{S}_{p}$ by the $\mathbb{Z}$-action generated by the map $\phi^{2}$. The conclusion follows from the fact that the square of a cyclic permutation of odd length is still cyclic, while the square of a cyclic permutation of even length factors as a product of two cycles of half the original length.
@@
In the following, it will be useful to consider the genus one Heegaard splitting of the three-sphere
@@
We can assume that the map $f_{p}$ in the proof of Lemma 2 is a rotation of $S^{3}$ of angle $2\pi/p$, which restricts to a rotation along the $D^{2}$-factor on the solid torus on the left side of (2.5) and to a rotation along the $S^{1}$-factor on the solid torus on the right side. Moreover, we can choose the three-balls $\{B_{1},\dots,B_{p}\}$ to be symmetrically embedded in the solid torus on the right side of (2.5), where they are permuted by the action of $f_{p}$.
@@
We denote by $\mathbb{A}:S^{2}\rightarrow S^{2}$ the antipodal involution.
[Proof]A result of Gluck [18] states that the diffeotopy group of $S^{1}\times S^{2}$ is isomorphic to $\mathbb{Z}/2\times\mathbb{Z}/2\times\mathbb{Z}/2$. This group is generated by the reflections $(\theta,x)\mapsto(\theta,\mathbb{A}(x))$ and $(\theta,x)\mapsto(-\theta,x)$ along with the self-diffeomorphism
@@
given by $(\theta,x)\mapsto(\theta,r_{\theta}(x))$, where $r_{\theta}:S^{2}\rightarrow S^{2}$ is a rotation of an angle $\theta$ around a fixed axis [18], [21, p. 156]. Such map is known as a Gluck twist. It is straight-forward to see that Theorem 4 holds for the first two generators of the diffeotopy group of $S^{1}\times S^{2}$. We now show that the diffeomorphism (2.6) admits an extension $\Phi_{p}:S_{p}\rightarrow S_{p}$ for any odd $p\in\mathbb{Z}_{>0}$. We explicitly build the diffeomorphism $\Phi_{p}$ by using the identification
(2.7) $S_{p}\cong(\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i}))/\sim_{\phi}$
coming from Lemma 2 as follows. Without loss of generality, we work in the setting of Remark 3: the deck transformation $\phi$ is defined in (2.3) and the collection of three-balls $\{B_{1},\dots,B_{p}\}\subset D^{2}\times S^{1}$ are defined as in the proof of Lemma 2.
@@
$F:\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})\rightarrow S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i}$
@@
where $r_{t}:S^{3}\rightarrow S^{3}$ is a rotation of an angle $2\pi t$ that restricts to a rotation along the $D^{2}$-factor on the solid torus $D^{2}\times S^{1}$ of the Heegaard decomposition of $S^{3}$ as described in Remark 3. This solid torus contains the three-balls $\{B_{1},\dots,B_{p}\}$. Notice that for every $t\in[0,1]$, the map $F(t,\cdot)$ is an $f_{p}$-equivariant self-diffeomorphism of $S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i}$, where $f_{p}$ can be chosen to act on the genus one Heegaard splitting of $S^{3}$ described in Remark 3. We can now define a self-diffeomorphism
(2.8) $\Phi:\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})\rightarrow\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})$
by
(2.9) $(t,x)\mapsto(t,F(t,x)).$
It is immediate to check that the diffeomorphism (2.8) is $\phi$-equivariant. Therefore, it induces a self-diffeomorphism $\Phi_{p}:S_{p}\rightarrow S_{p}$. In particular, the restriction of $\Phi_{p}$ to $\partial S_{p}=S^{1}\times S^{2}$ is a composition of $p$ Gluck twists, which is again a Gluck twist since $p$ is odd. This concludes the proof.
Our construction and Theorem 4 also hold in the topological category. The reader is directed to [17, Chapter 5] regarding existence and uniqueness results of the tubular neighborhood of a locally flat simple loop in a topological four-manifold. As mentioned in the first step of our introduction, remove the tubular neighborhood of a locally flat framed simple loop $\alpha_{M}\subset M$ from a closed oriented topological four-manifold $M$ to obtain a compact oriented topological four-manifold $M\setminus\nu(\alpha_{M})$ with boundary $\partial(M\setminus\nu(\alpha_{M}))=S^{1}\times S^{2}$. Cap off its boundary with the compact topological four-manifold defined in (2.1) to construct a closed oriented topological four-manifold
@@
[Proof]We prove the third item of the lemma. Let $\alpha_{M}\subset M$ be a simple loop as in (2.10) and denote by $\widetilde{\alpha}_{M}\subset\widetilde{M}$ its preimage in the universal cover $\widetilde{M}$ of $M$. The universal cover $\widetilde{M}(p)$ of $M(p)$ is constructed by gluing together copies of $\widetilde{M}\setminus\nu(\widetilde{\alpha}_{M})$ and $\widetilde{S}_{p}/\sim_{\phi^{k}}$ by matching each boundary component of $\widetilde{M}\setminus\nu(\widetilde{\alpha}_{M})$ with a boundary component of $\widetilde{S}_{p}/\sim_{\phi^{k}}$, where $\widetilde{S}_{p}$ is the universal cover of $S_{p}$ (see Lemma 2) and $k$ is the order of the homotopy class $\alpha_{M}]\in\pi_{1}(M)$. In particular, if $k=\infty$ we set $\widetilde{S}_{p}/\sim_{\phi^{\infty}}=\widetilde{S}_{p}$. It follows that the equivariant intersection form $\widetilde{I}_{M(p)}$ of $M(p)$ is
@@
A proof of the other items of Lemma 6 is obtained from straight-forward homological computations, Novikov’s additivity [21, Remark 9.1.7] and the Seifert-van Kampen theorem.
@@
Let $M$ be a closed oriented four-manifold and let $\alpha_{M}\subset M$ be a simple framed loop whose homotopy class is $\alpha_{M}]=0\in\pi_{1}(M)$. There is a diffeomorphism